# Oze: Decentralized Graph-based Concurrency Control for Long-running Update Transactions (Extended Version)

### v2

@@ 2.6. BoMB with Conventional Protocols and Challenges
Figure 4 illustrates why existing concurrency control protocols fail to handle BoMB transactions. Inherently, transaction serializability can be determined solely by the dependency order, i.e., reads-from relation and version order, including anti-dependencies decided by version order (Weikum and Vossen, 2001). To reduce the cost of tracking all of those dependencies, conventional protocols impose a certain constraint on the acceptable serialization order. This protocol-specific constraint is an order, and we call it constraint order. If the observed dependency order (including potential ones) of a transaction does not match the constraint order, the transaction is aborted. Most of the existing protocols abort the L1 transaction because the dependency order of BoMB transactions (S2$<$L1$<$S1) does not obey their own constraint orders, although this is a false positive.
For example, OCC protocols such as Silo (Tu et al., 2013) and TicToc (Yu et al., 2016) force transactions to obey a constraint order, which is an order in which they entered the serialization point by acquiring the necessary locks in the validation phase. The OCC-specific behavior, such as aborts due to finding an overwriter, can be explained as a mismatch between a potential anti-dependency order (L1$<$S1) and the constraint order (S1$<$L1), as shown in Figure 4(a). Due to the constraint order, the L1 transaction is aborted as a false positive even though they are serializable (S2$<$L1$<$S1).
@@
MVTO forces transactions to obey the begin timestamp order as the constraint order. Because the observed anti-dependency (S2$<$L1) of an L1 transaction is opposite to the begin timestamp order (L1$<$S2), the L1 transaction is aborted, as shown in Figure 4(b).
@@ 4.3.4. Complexity
[…]
@@ 5.1. Experimental Setup
(a) L1 length scalability with L1 abort rate <1<1< 1% (1 thread for L1 and 1 thread for other short transactions)
(b) Thread scalability with 100 products and L1 abort rate <1<1< 1% (1 thread for L1 and varying # threads for short transactions)
@@
(a) L1 length scalability with L1 abort rate <1<1< 1% (1 thread for L1 and 1 thread for other short transactions)
(b) Thread scalability with 100 products and L1 abort rate <1<1< 1% (1 thread for L1 and varying # threads for short transactions)

### v3

@@ 2.6. BoMB with Conventional Protocols and Challenges
Figure 4 illustrates why existing concurrency control protocols fail to handle BoMB transactions. Inherently, transaction serializability can be determined solely by the dependency order, i.e., reads-from relation and version order, including anti-dependencies decided by version order (Weikum and Vossen, 2001). To reduce the cost of tracking all of those dependencies, conventional protocols impose a certain constraint on the acceptable serialization order. This protocol-specific constraint is an order, and we call it constraint order. If the observed dependency order (including potential ones) of a transaction does not match the constraint order, the transaction is aborted. Most of the existing protocols abort the L1 transaction because the dependency order of BoMB transactions (S2<<L1<<S1) does not obey their own constraint orders, although this is a false positive.
For example, OCC protocols such as Silo (Tu et al., 2013) and TicToc (Yu et al., 2016) force transactions to obey a constraint order, which is an order in which they entered the serialization point by acquiring the necessary locks in the validation phase. The OCC-specific behavior, such as aborts due to finding an overwriter, can be explained as a mismatch between a potential anti-dependency order (L1<<S1) and the constraint order (S1<<L1), as shown in Figure 4(a). Due to the constraint order, the L1 transaction is aborted as a false positive even though they are serializable (S2<<L1<<S1).
@@
MVTO forces transactions to obey the begin timestamp order as the constraint order. Because the observed anti-dependency (S2<<L1) of an L1 transaction is opposite to the begin timestamp order (L1<<S2), the L1 transaction is aborted, as shown in Figure 4(b).
@@ 4.3.4. Complexity
The dominant factor of the complexity in the validation phase is graph processing, such as merging and cycle-checking. The time complexities of both processes are $O(|V|+|E|)$ where $|V|$ and $|E|$ are the number of nodes and edges in a graph [^2] Precisely, the cycle check requires fewer nodes and edges since it is enough to only check the nodes that follow the transaction about to commit., respectively. The graph processing must be done for each record in the target record set $X_{i}$. Therefore, as a whole, the time complexity of the decentralized Oze protocol is $O(|X_{i}|(|V|+|E|))$.
@@ 5.1. Experimental Setup
(a) L1 length scalability with L1 abort rate <1<1% (1 thread for L1 and 1 thread for other short transactions)
(b) Thread scalability with 100 products and L1 abort rate <1<1% (1 thread for L1 and varying # threads for short transactions)
@@
(a) L1 length scalability with L1 abort rate <1<1% (1 thread for L1 and 1 thread for other short transactions)
(b) Thread scalability with 100 products and L1 abort rate <1<1% (1 thread for L1 and varying # threads for short transactions)